solve the quadratic equation 2xpower 2 +ax-a power 2=0 for x.

Grade:12th pass

1 Answers

Arun Kumar IIT Delhi
9 years ago
Hello Student,

2x^{2}+ax-a^{2}=0
2x^{2}+2ax-ax-a^{2}=0
2x(x+a)-a(x+a)=0
(2x-a)(x+a)=0
x = a/2, -a

Thanks & Regards
Arun Kumar
Btech, IIT Delhi
Askiitians Faculty